OpenStudy (anonymous):

I need help /: 3(x + 1) = −2(x − 1) − 4

OpenStudy (anonymous):

3x + 3 = -2x + 2 - 4 5x = 2 - 4 - 3 5x = -5 x = -1

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

;)